OpenStudy (anonymous):

what is the product of -3x^2y and (5xy^2+xy)

OpenStudy (anonymous):

-15x^(y+1)y^2 -3x^(y+1)*y

OpenStudy (anonymous):

-3 y (5 y+1) x^(y+1)

OpenStudy (anonymous):

-15x^y+1*y^2-3x^y+1.y

OpenStudy (whpalmer4):

\[-3x^2y*(5xy^2+xy) = -3*5*x^2*x*y*y^2 -3*x^2*x*y*y \]\[= -15x^3y^3-3x^3y^2\]which can be factored as\[-3x^3y^2(5y+1)\]